# primordial black holes; microlensing; dark matter

### v1

@@ Abstract
Stellar microlensing is a powerful tool for probing dark matter in the form of planetary and stellar mass compact objects (COs), in particular primordial black holes (PBHs). Under standard assumptions, current observations exclude COs in the mass range $10^{-11}\lesssim M/M_{\odot}\lesssim 10^{4}$aking up all of the dark matter. We provide an overview, aimed at theorists working on PBHs, of the history, theory, observational status, and future prospects of the field.
@@ 1 Introduction and history
In 2016 the discovery of gravitational waves from mergers of multi-Solar mass PBHs by LIGO-Virgo Abbott et al. (2016) generated a surge of interest in PBH dark matter (DM) Bird et al. (2016); Clesse and García-Bellido (2017); Sasaki et al. (2016); Carr et al. (2016). However microlensing observations over two decades by the OGLE collaboration have failed to find the large, $\gtrsim 10^{2}$, number of events expected (under standard assumptions) if a significant fraction of the MW halo is in the form of stellar mass CO Mróz et al. (2024). These long duration observations have also increased the largest CO mass probed to $\sim 10^{4}M_{\odot}$, while high cadence observations by Niikura et al. have decreased the lowest mass to $\sim 10^{-11}M_{\odot}$ Niikura et al. (2019).
We overview how stellar microlensing surveys probe the fraction of COs, such as PBHs, in the MW halo by observing stars in the MC or M31. Our aim is to provide an introduction to this field that is accessible to theorists who are interested PBHs, and specifically the possibility that planetary and stellar mass PBHs make up a significant fraction of the DM. First, in Sec. 2, we review the theory of stellar microlensing, before focusing on observations in Sec. 3. We then look at future prospects in Sec. 4, before concluding with a brief summary in Sec. 5.
@@ 2.1.1 Individual events
[…]
Figure 1: The geometry of a stellar microlensing event. The left and right panels show the view perpendicular to and along the line of sight respectively. The solid black and dashed grey lines shows the trajectory of the compact object and the Einstein radius, $R_{\rm E}$, respectively. The orange star, black circle and blue square denote the source, the lens and the observer respectively. In the left panel the dotted green lines show the distance from the observer to the lens and source, $D_{\rm L}=xD_{\rm S}$ and $D_{\rm S}$ respectively. In the right panel the green dotted lines show the impact parameter, $b$, and its minimum value $b_{\rm min}$.
@@
$\hat{t}\sim 1\,{\rm yr}\,\sqrt{x(1-x)}\left(\frac{M}{M_{\odot}}\right)^{1/2}\,.$ (4)
Due to the mass dependence of the timescale, observations over a long duration are need to probe large masses, while high cadence observations (i.e. at short intervals) are required to probe small masses.
@@
where $u_{\rm min}=b_{\rm min}/R_{\rm E}$, $b_{\rm min}$ is the minimum value of the distance of the lens from the line of sight to the source, and time $t$ is defined so that $t=0$ when $b=b_{\rm min}$. Figure 1 shows the geometry of a microlensing event, as viewed parallel to and along the line of sight, including $R_{\rm E}$, $D_{\rm S}$, $D_{\rm L}$, $b$ and $b_{\rm min}$. Figure 2 shows the magnification factor $A$, given by Eq. (5), as a function of $2t/\hat{t}$ for different values of $u_{\rm min}$. This light curve, which is symmetric and achromatic (i.e. the same for all wavelengths), is often referred to as the Paczyński light curve.
@@ 2.1.2 Optical depth and event rate
A more useful quantity is the rate, $\Gamma$}, at which events occur, and we summarize the calculation as outlined in Ref. Mao (2012). The area, ${\rm d}\tilde{A}$, swept out by a lens in a time interval, ${\rm d}t$, is given by the product of the diameter of the Einstein ring, $2R_{\rm E}$, and the distance traveled, $v_{\rm t}{\rm d}t$: $^{\rm d}\tilde{A}=2R_{\rm E}v_{\rm t}{\rm d}t=4R_{\rm E}^{2}{\rm d}t/\hat{t}$. The probability, ${\rm d}\tau$, of an individual star undergoing a new microlensing event is given by
${\rm d}\tau=\int_{0}^{D_{\rm S}}n(D_{\rm L})\,{\rm d}A\,{\rm d}D_{\rm L}=\int_{0}^{D_{\rm S}}n(D_{\rm L})\left(\frac{4R_{\rm E}^{2}}{\hat{t}}\right)\,{\rm d}t\,{\rm d}D_{\rm L}\,.$ (7)
The total number of new microlensing events in time ${\rm d}t$ is $N_{\star}{\rm d}\tau$, where $N_{\star}$ is the number of stars being monitored, and hence the event rate is
$\Gamma=N_{\star}\frac{{\rm d}\tau}{{\rm d}t}=N_{\star}\int_{0}^{D_{\rm S}}n(D_{\rm d})\left(\frac{4}{\pi\hat{t}}\right)\pi R_{\rm E}^{2}{\rm d}D_{\rm d}=\frac{4N_{\star}}{\pi}\int\frac{{\rm d}\tau}{\hat{t}}\,$. (8)
@@ 2.2 Differential event rate and parameter constraints
however this assumption is only exact if the density varies with Galactocentric radius $r$ as $\rho(r)\propto r^{-2}$ for all $r$.
@@
The standard halo model historically assumed in microlensing studies (‘Model S’ in e.g. Ref. Alcock et al. (1997)) is a cored isothermal sphere with a density profile
@@
Sahu pointed out that stars ($\star$’) in the LMC can act as lenses Sahu (1994), a phenomenon often referred to as ’self-lensing’. There are also contributions to the total differential event rate from stars in the MW disc and and COs in the LMC:
@@
The self-lensing rate is largest for observing fields in the centre of the LMC, where the number of stars is highest. The MACHO collaboration found, using models from Ref. Gyuk et al. (2000), an optical depth $\tau_{{\rm LMC},\star}\sim(1-3)\times 10^{-8}$. Using the same LMC model with some minor modifications, OGLE III and IV (which has a wider observing field) found 5.7 events expected from stars in the LMC (compared with the $10^{2-3}$ expected if all of the halo is in planetary or stellar mass COs) Mróz et al. (2024). The MACHO collaboration found an optical depth for microlensing by stars in the MW disc, $\tau_{{\rm MW},\star}\lesssim 10^{-8}$, lower than their estimate for self-lensing Alcock et al. (2000). However more recent studies have found significantly larger MW disc rates, comparable with, or even larger, than the self-lensing rate, in part due to a larger thick disc contribution Calchi Novati and Mancini (2011); Mróz et al. (2024). For the two MW models they considered Han and Gould (2003); Cautun et al. (2020), OGLE III and IV found 7.0 and 14.7 expected events respectively from stars in the MW disc Mróz et al. (2024). Note, however, that the relative numbers of events from different lens populations depend on the fields observed, and the detection efficiency (see below).
@@
where $N_{\rm obs}$ is the number of observed events and $\hat{t}_{i}$ are their timescales.
The MACHO first year results paper Alcock et al. (1997) studied a very wide range of models for the MW halo, some of which are not consistent with current observations. They found that there was a significant variation in the differential event rate, and hence the observational constraints on the CO mass and halo fraction. The OGLE collaboration finds that the uncertainty in their constraint is relatively small (see Extended Data Fig. 1 of Ref. Mróz et al. (2024)), however they only consider two benchmark models (with fixed parameter values) and assume a Maxwellian velocity distribution. Other studies Hawkins (2015); Green (2017); Calcino et al. (2018); Garcia-Bellido and Hawkins (2024) have found that the uncertainty in exclusion limits can be significant. Ref. Green (2025) found that the rate of long-timescale events (and hence the constraints on low-mass WIMPs) has a fairly weak dependence on the CO density and velocity distribution. However the rate of short-timescale events (and hence the constraints on high-mass WIMPs) depends sensitively on the local DM velocity distribution. An accurate calculate of the differential event rate (and parameter constraints) for models consistent with current observations, and numerical simulations of Milky Way-like galaxies, is an outstanding issue.
@@
The standard calculation of the halo differential event rate outlined above, Eq. (10), assumes that the dark matter distribution is smooth, and fully specified by the density profile, $\rho(r)$. However PBH dark matter is expected to cluster more on sub-galactic scales than particle dark matter. Since PBHs are discrete objects, there are Poisson fluctuations in their distribution and PBH clusters start forming shortly after matter-radiation equality Afshordi et al. (2003); Inman and Ali-Haïmoud (2019). For PBHs formed from gaussian fluctuations, the clusters are sufficiently diffuse Jedamzik (2020) that the PBHs still act as individual microlenses. In this case the effect on the microlensing differential event rate, and hence constraints on the PBH mass and halo fraction, is small Petač et al. (2022); Gorton and Green (2022). Some PBHs are expected to be in binaries which could potentially lead to a source being lensed twice, however the binaries are typically sufficiently wide that the time between such repeating events is expected to be much longer that the duration of microlensing surveys Petač et al. (2022). The probability distribution of perturbations that are large enough to produce PBHs which make up a significant fraction of the dark matter, are expected to be non-gaussian, e.g. Ref. Atal and Germani (2019), however, and PBHs formed from non-gaussian fluctuations can form more compact clusters Young and Byrnes (2020). If clusters are sufficiently compact that the cluster as a whole acts as a microlens, then microlensing parameter constraints (or allowed regions) will shift to lower PBH masses Calcino et al. (2018).
@@ 2.3 Non-standard events
The standard light curve is significantly modified if the impact parameter of an event is comparable to the ratio of the radius of the source star to the Einstein radius Witt and Mao (1994); Gould (1994). In particular finite source effects significantly reduce the maximum magnification for compact objects with mass $M\lesssim 10^{-9}M_{\odot}$ Sasaki et al. (2018).
@@
PBHs with mass $M_{\rm PBH}\lesssim 10^{-11}M_{\odot}$ have Schwarzschild radius comparable to the wavelength of light, resulting in diffraction and interference effects. These wave optics effects reduce the magnification and place a lower limit on the mass of COs that can be probed by microlensing Sugiyama et al. (2020); Montero-Camacho et al. (2019).
@@ 2.4 Astrometric microlensing
$\pi_{\rm rel}=\pi_{\rm L}-\pi_{\rm S}=\frac{1}{D_{\rm L}}-\frac{1}{D_{\rm S}}=\frac{D_{\rm S}-D_{\rm L}}{D_{\rm L}D_{\rm S}}\,$, (18)
in milliarcseconds (mas), as
$\theta_{\rm E}^{2}=8.144\,{\rm mas}\left(\frac{M}{M_{\odot}}\right)\left(\frac{\pi_{\rm rel}}{{\rm mas}}\right)\,.$ (19)
@@ 3 Observations
In this Section we first overview the general principles of stellar microlensing observations (Sec. 3.1), before looking at the results of specific surveys which have published constraints on the halo fraction of COs in Secs. 3.2-3.6. Unless otherwise mentioned, all limits (or parameter values) quoted assume the standard halo model, see Sec. 2.2. Sec. 3.7 contains a summary of the current observational status. Figure 4 shows the final published parameter constraints on the halo fraction, $f$, and CO mass, $M$, assuming all CO have the same mass. Table 2 summarizes the main features of the key LMC microlensing surveys, including the number of events they observed. This Section aims to provide theorists with an introduction to the general principles and current status of microlensing surveys. For further details see Ref. Moniez (2010), a 2010 review of microlensing from an observational perspective.
@@ 3.2 MACHO
As part of their response to criticism by Hawkins and García-Bellido Hawkins and García-Bellido (2025), the OGLE collaboration have published a reanalysis of the MACHO candidate events using more modern DIA techniques Mróz et al. (2025). They deem events MACHO-LMC-6, 13 and 15 weak candidates, due to their light curves being either asymmetric, or lacking sufficient data around the peak. They note, however, that Bennett et al. Bennett et al. (2005) presented additional follow-up photometry which supported the microlensing interpretation of events 13 and 15. Based on their new DIA light curves, they conclude that 7 of the MACHO events are very likely microlensing (MACHO-LMC-1, 4, 5, 9, 14, 21, 15) and argue that the other candidate events would be unlikely to pass ’sensible microlensing cuts’.
Table  1 contains a summary of the current status of the 13 events which passed criteria A. Of the 13 events, 4 are now thought to be variable stars and 2 are microlensing events with the lens likely not in the MW halo. This leaves 6 or 7 candidate lensing events where the lens could be located in the MW halo.
@@
Table 1: A summary of the current status of the 13 events which passed criteria A in the 5.7 year MACHO collaboration analysis Alcock et al. (2000). See also Table 1 of Ref. Mróz et al. (2025).
@@
Event Status
1 lens is possibly binary Dominik and Hirshfeld (1994, 1996); Mróz et al. (2025)
@@
5 lens is star in MW disk Drake et al. (2004)
6
7 variable star (has varied repeatedly) Wyrzykowski et al. (2011)
8 more likely a variable star Mróz et al. (2024)
13
14 lens likely in LMC Alcock et al. (2001)
15
18 more likely a variable star Mróz et al. (2024)
@@
23 brightened a 2nd time, likely a variable star Bennett et al. (2005)

### v2

@@ Abstract
Stellar microlensing surveys are a powerful tool for probing dark matter in the form of planetary and stellar mass compact objects (COs), in particular primordial black holes (PBHs). Under standard assumptions, current observations exclude COs in the mass range $10^{-11}\lesssim M/M_{\odot}\lesssim 10^{4}$ making up all of the dark matter. We provide an overview, aimed at theorists working on PBHs, of the history, theory, observational status, and future prospects of the field.
@@ 1 Introduction and history
In 2016 the discovery of gravitational waves from mergers of multi-Solar mass PBHs by LIGO-Virgo Abbott et al. (2016) generated a surge of interest in PBH dark matter (DM) Bird et al. (2016); Clesse and García-Bellido (2017); Sasaki et al. (2016); Carr et al. (2016). However microlensing observations over two decades by the OGLE collaboration have failed to find the large, $\gtrsim 10^{2}$, number of events expected (under standard assumptions) if a significant fraction of the MW halo is in the form of stellar mass COs Mróz et al. (2024). These long duration observations have also increased the largest CO mass probed to $\sim 10^{4}M_{\odot}$, while high cadence observations by Niikura et al. have decreased the lowest mass to $\sim 10^{-11}M_{\odot}$ Niikura et al. (2019).
We overview how stellar microlensing surveys probe the fraction of COs, such as PBHs  [^1] Throughout, we use CO when covering material that applies to compact objects in general, and PBH for material that is specific to Primordial Black Holes., in the MW halo by observing stars in the MC or M31  [^2] We do not cover constraints on CO DM Diego et al. (2018); Müller and Miralda-Escudé (2024) from supermagnified stars (a star in a galaxy that is magnified by a large factor by lensing by a galaxy cluster, with the star then magnified further by microlensing by a CO) Venumadhav et al. (2017).. Our aim is to provide an introduction to this field that is accessible to theorists who are interested in PBHs, and specifically the possibility that planetary and stellar mass PBHs make up a significant fraction of the DM. First, in Sec. 2, we review the theory of stellar microlensing, before focusing on observations in Sec. 3. We then look at future prospects in Sec. 4, before concluding with a brief summary in Sec. 5.
@@ 2.1.1 Individual events
$\hat{t}\sim 1\,{\rm yr}\,\sqrt{x(1-x)}\left(\frac{M}{M_{\odot}}\right)^{1/2}\left(\frac{v_{\rm t}}{200\,{\rm km\,s}^{-1}}\right)^{-1}\,.$ (4)
Due to the mass dependence of the timescale, observations over a long duration are needed to probe large masses, while high cadence observations (i.e. at short intervals) are required to probe small masses.
@@
[…]
Figure 1: The geometry of a stellar microlensing event. The left and right panels show the view perpendicular to and along the line of sight respectively. The solid black and dashed grey lines show the trajectory of the compact object and the Einstein radius, $R_{\rm E}$, respectively. The orange star, black circle and blue square denote the source, the lens and the observer respectively. In the left panel the dotted green lines show the distance from the observer to the lens and source, $D_{\rm L}=xD_{\rm S}$ and $D_{\rm S}$ respectively. In the right panel the green dotted lines show the impact parameter, $b$, and its minimum value $b_{\rm min}$.
@@
where $u_{\rm min}=b_{\rm min}/R_{\rm E}$, $b_{\rm min}$ is the minimum value of the distance of the lens from the line of sight to the source, and time $t$ is defined so that $t=0$ when $b=b_{\rm min}$. Figure 1 shows the geometry of a microlensing event, as viewed perpendicular to and along the line of sight, including $R_{\rm E}$, $D_{\rm S}$, $D_{\rm L}$, $b$ and $b_{\rm min}$. Figure 2 shows the magnification factor $A$, given by Eq. (5), as a function of $2t/\hat{t}$ for different values of $u_{\rm min}$. This light curve, which is symmetric and achromatic (i.e. the same for all wavelengths), is often referred to as the Paczyński light curve.
@@ 2.1.2 Optical depth and event rate
A more useful quantity is the rate, $\Gamma$}, at which events occur, and we present a modified version of the illustrative calculation outlined in Ref. Mao (2012). The area, ${\rm d}\tilde{A}$, swept out by a lens in a time interval, ${\rm d}t$, is given by the product of the diameter of the Einstein ring, $2R_{\rm E}$, and the distance traveled, $v_{\rm t}{\rm d}t$: $^{\rm d}\tilde{A}=2R_{\rm E}v_{\rm t}{\rm d}t=4R_{\rm E}^{2}{\rm d}t/\hat{t}$. The probability, ${\rm d}\tau$, of an individual star undergoing a new microlensing event is given by
${\rm d}\tau=n(D_{\rm L})\,\pi R_{\rm E}^{2}\,{\rm d}D_{\rm L}\,.$ (7)
If $N_{\star}$ is the number of stars being monitored, then the number of new microlensing events in time ${\rm d}t$, ${\rm d}N$, is given by
${\rm d}N=N_{\star}\int_{0}^{D_{\rm S}}n(D_{\rm d}){\rm d}\tilde{A}\,{\rm d}D_{\rm d}=N_{\star}\int_{0}^{D_{\rm S}}n(D_{\rm d})\frac{4R_{\rm E}^{2}}{\hat{t}}\,{\rm d}t\,{\rm d}D_{\rm d}=\frac{4N_{\star}}{\pi}\int_{0}^{D_{\rm S}}\frac{1}{\hat{t}}{\rm d}\tau\,{\rm d}t\,$, (8)
and hence the rate at which new events occur, $\Gamma$, is given by
$\Gamma\equiv\frac{{\rm d}N}{{\rm d}t}=\frac{4N_{\star}}{\pi}\int_{0}^{D_{\rm S}}\frac{{\rm d}\tau}{\hat{t}}\,$. (9)
@@ 2.2 Differential event rate and parameter constraints
however this is only true for an isotropic, isothermal sphere.
@@
The standard halo model historically assumed in microlensing studies (‘Model S’ in e.g. Ref. Alcock et al. (1997)) has a density profile
@@
There are also contributions to the total differential event rate, $^{\rm d}\Gamma/{\rm d}\hat{t}$, from stars ($\star$’) in the MW disc and LMC, and COs in the LMC:
@@
Lensing by stars in the LMC, a possibility that was pointed out by Sahu Sahu (1994), is often referred to as ’self-lensing’. The self-lensing rate is largest for observing fields in the centre of the LMC, where the number of stars is highest. The MACHO collaboration found, using models of the LMC disc from Ref. Gyuk et al. (2000), an optical depth $\tau_{{\rm LMC},\star}\sim(1-3)\times 10^{-8}$. Using the same LMC model with some minor modifications, OGLE III and IV (which has a wider observing field) found 5.7 events expected from stars in the LMC (compared with the $10^{2-3}$ expected if all of the halo is in planetary or stellar mass COs) Mróz et al. (2024). The MACHO collaboration found an optical depth for microlensing by stars in the MW disc, $\tau_{{\rm MW},\star}\lesssim 10^{-8}$, lower than their estimate for self-lensing Alcock et al. (2000). However more recent studies have found significantly larger MW disc rates, comparable with, or even larger, than the self-lensing rate, in part due to a larger thick disc contribution Calchi Novati and Mancini (2011); Mróz et al. (2024). For the two MW models they considered Han and Gould (2003); Cautun et al. (2020), OGLE III and IV found 7.0 and 14.7 expected events respectively from stars in the MW disc Mróz et al. (2024). Note, however, that the relative numbers of events from different lens populations depend on the fields observed, and the detection efficiency (see below).
@@
where $N_{\rm obs}$ is the number of observed events and $\hat{t}_{i}$ are their timescales. PBHs are expected to have an extended mass function (see e.g. Sec. 2 of Ref. Carr et al. (2026)). Constraints calculated assuming all CO have the same mass can be translated into constraints for a specific extended mass function using the method presented in Ref. Carr et al. (2017). Furthermore, as pointed out in Ref. Alcock et al. (1996), if $f=1$ is excluded for a particular range of masses under the assumption that all CO have the same mass, then any extended mass function with $f=1$ in this range is excluded. We note that the constraints apply to any COs in the MW halo, however (as mentioned in the introduction) astrophysical objects, such as white dwarfs, are excluded from making up a large fraction of the halo by baryon budget arguments Fields et al. (2000). Ref. Croon et al. (2020) has recalculated the microlensing constraints for extended objects, such as boson stars.
The MACHO first year results paper Alcock et al. (1997) studied a very wide range of models for the MW halo, some of which are not consistent with current observations. They found that there was a significant variation in the differential event rate, and hence the observational constraints on the CO mass and halo fraction. The OGLE collaboration finds that the uncertainty in their constraint is relatively small (see Extended Data Fig. 1 of Ref. Mróz et al. (2024)), however they only consider two benchmark models (with fixed parameter values) and assume a Maxwellian velocity distribution. Other studies Hawkins (2015); Green (2017); Calcino et al. (2018); Garcia-Bellido and Hawkins (2024) have found that the uncertainty in exclusion limits can be significant. Ref. Green (2025) found that the rate of long-timescale events (and hence the constraints on low-mass COs) has a fairly weak dependence on the CO density and velocity distribution. However the rate of short-timescale events (and hence the constraints on high-mass COs) depends sensitively on the local DM velocity distribution. An accurate calculation of the uncertainties on the differential event rate (and parameter constraints) for models consistent with current observations, and numerical simulations of Milky Way-like galaxies, is an outstanding issue.
@@
Microlensing surveys towards the Galactic bulge can be used to probe CO DM (see Secs. 3.4 and 4.1). However in this case it is necessary to distinguish events due to CO DM in the MW halo from the large population of events caused by astrophysical objects in the luminous components of the MW.
@@
The standard calculation of the halo differential event rate outlined above, Eq. (11), assumes that the CO dark matter distribution is smooth, and fully specified by the density profile, $\rho(r)$. However PBH dark matter is expected to cluster more on sub-galactic scales than particle dark matter. Since PBHs are discrete objects, there are Poisson fluctuations in their distribution and PBH clusters start forming shortly after matter-radiation equality Afshordi et al. (2003); Inman and Ali-Haïmoud (2019). For PBHs formed from gaussian fluctuations, the clusters are sufficiently diffuse Jedamzik (2020) that the PBHs still act as individual microlenses. In this case the effect on the microlensing differential event rate, and hence constraints on the PBH mass and halo fraction, is small Petač et al. (2022); Gorton and Green (2022). Some PBHs are expected to be in binaries which could potentially lead to a source being lensed twice, however the binaries are typically sufficiently wide that the time between such repeating events is expected to be much longer that the duration of microlensing surveys Petač et al. (2022). The probability distribution of perturbations that are large enough to produce PBHs which make up a significant fraction of the dark matter, are expected to be non-gaussian, e.g. Ref. Atal and Germani (2019), however, and PBHs formed from non-gaussian fluctuations can form more compact clusters Young and Byrnes (2020). If clusters are sufficiently compact that the cluster as a whole acts as a microlens, then microlensing parameter constraints (or allowed regions) will shift to lower PBH masses Calcino et al. (2018).
@@ 2.3 Non-standard events
The standard light curve is significantly modified if the impact parameter of an event is comparable to the ratio of the radius of the source star to the Einstein radius Witt and Mao (1994); Gould (1994). In particular finite source effects significantly reduce the maximum magnification for COs with mass $M\lesssim 10^{-9}M_{\odot}$ Sasaki et al. (2018).
@@
PBHs with mass $M_{\rm PBH}\lesssim 10^{-11}M_{\odot}$ have Schwarzschild radius comparable to the wavelength of light, resulting in diffraction and interference effects. These wave optics effects reduce the magnification and place a lower limit on the mass of PBHs that can be probed by microlensing Sugiyama et al. (2020); Montero-Camacho et al. (2019).
@@ 2.4 Astrometric microlensing
$\frac{\pi_{\rm rel}}{\rm mas}=\frac{\pi_{\rm L}}{\rm mas}-\frac{\pi_{\rm S}}{\rm mas}=\frac{\rm kpc}{D_{\rm L}}-\frac{\rm kpc}{D_{\rm S}}\,$, (19)
in milliarc (mas), as
$\theta_{\rm E}^{2}=8.144\,{\rm mas}\left(\frac{M}{M_{\odot}}\right)\pi_{\rm rel}\,.$ (20)
@@ 3 Observations
In this Section we first overview the general principles of stellar microlensing observations (Sec. 3.1), before looking at the results of specific surveys which have published constraints on the halo fraction of COs in Secs. 3.2-3.6. Unless otherwise mentioned, all limits (or parameter values) quoted assume that the MW halo is described by the standard halo model, see Sec. 2.2, and that all CO have the same mass. Sec. 3.7 contains a summary of the current observational status. Figure 4 shows the final published parameter constraints on the halo fraction, $f$, and CO mass, $M$. Table 2 summarizes the main features of the key LMC microlensing surveys, including the number of events they observed. This Section aims to provide theorists with an introduction to the general principles and current status of microlensing surveys. For further details see Ref. Moniez (2010), a 2010 review of microlensing from an observational perspective.
@@ 3.2 MACHO
As part of their response to criticism by Hawkins and García-Bellido Hawkins and García-Bellido (2025), the OGLE collaboration have published a reanalysis of the MACHO candidate events using more modern DIA techniques Mróz et al. (2025). They deem events MACHO-LMC-6, 13 and 15 weak candidates, due to their light curves being either asymmetric, or lacking sufficient data around the peak. They note, however, that Bennett et al. Bennett et al. (2005) presented additional follow-up photometry which supported the microlensing interpretation of events 13 and 15. Based on their new DIA light curves, they conclude that 7 of the MACHO events are very likely microlensing (MACHO-LMC-1, 4, 5, 9, 14, 21, 25) and argue that the other candidate events would be unlikely to pass ’sensible microlensing cuts’.
Table  1 contains a summary of the current status of the 13 events which passed criteria A. Of the 13 events, 4 are now thought to be variable stars and 2 are microlensing events with the lens likely not in the MW halo. This leaves at most 7 candidate lensing events where the lens could be located in the MW halo.
@@
Table 1: A summary of the current status of the 13 events which passed criteria A in the 5.7 year MACHO collaboration analysis Alcock et al. (2000) (see also Table 1 of Ref. Mróz et al. (2025)). The 2nd column (’Microlensing’) contains information on whether the event is currently thought to be microlensing. Events classified as ’very likely microlensing’ and ’weak candidate’ by Ref. Mróz et al. (2025) are denoted by a green tick, , and an orange question mark, ?, respectively. Candidate events that are now thought to be variable stars are denoted by a red cross, . The 3rd column (’Lens’) contains information about the lens. It is left blank where there is no information and contains ’n/a’ for events which are now thought to be variable stars.
@@
Event Microlensing Lens
1 possibly binary Dominik and Hirshfeld (1994, 1996); Mróz et al. (2025)
@@
5 star in MW disk Drake et al. (2004)
6 ?
7 variable star (has varied repeatedly) Wyrzykowski et al. (2011) n/a
8 likely a variable star Mróz et al. (2024) n/a
13 ?
14 likely in LMC Alcock et al. (2001)
15 ?
18 likely a variable star Mróz et al. (2024) n/a
@@
23 brightened a 2nd time, likely a variable star Bennett et al. (2005) n/a